# Fast and Tunable Decoherence-Free Subspace Engineering

### v1

@@ I Introduction
A decoherence-free subspace (DFS) allows for exactly this, by taking advantage of the structure of quantum decoherence in order to remove it completely in a quantum system [24, 25, 26, 27, 28]. In a DFS, one engineers the dissipation to stabilize the desired quantum states against the dominant external noise sources. A promising approach in engineering dissipation is the use of dissipative cavity modes, where one controls the coupling of an atomic ensemble to the cavity mode via a dynamical driving protocol [29, 27, 30, 31]. These protocols have the advantage of removing the need for full control over all constituents of the quantum system, while the global coupling to the cavity mode creates a collective DFS. Such implementations have been discussed in the context of quantum computation and quantum sensing [32, 33, 34, 35, 36] where they allow stable coherence over longer timescales. However, the reduction of the required control in such implementations comes at the cost of their universality because, in general, they only protect against a subset of all possible errors. As a result, useful platforms and protocols which formulate a highly controllable DFS are extremely valuable.
@@
In particular, states can be prepared using time dependent dynamics where the state remains in an instantaneous DFS at every point in time. This allows states to be prepared using adiabatic shortcuts [45, 24] employing experimentally relevant driving schemes. Within this system, we show that a transitionless quantum driving method is possible. These methods allow one to prepare a decoherence-free state without any loss of purity and in timescales not limited by any error rate. We contrast this with typical methods of preparing a DFS which require dissipative state preparation [36] or adiabatic change to the system [46, 47, 48], and take orders of magnitude longer than our adiabatic shortcuts. These shorter timescales mean there is far less opportunity to have errors against which the DFS does not protect. This provides a step towards realizing a controllable DFS for quantum simulation [29, 49] and sensing [50, 51, 52, 53, 54] in atom-cavity setups.
This article is organized as follows. Sec. II is an outline of the basic requirements for a subspace to be a stable DFS. In Sec. III we introduce the collective atom-cavity model which will be used to create DFSs. From this, in Sec. IV we calculate the eigenstates of the decoherence operator to determine the dimension and structure of every DFS in the system. Next, Sec. V is a discussion of various time evolution methods used to prepare the system in a DFS. In Sec. VI we conclude with a discussion of future work.
@@ II Stable Decoherence-Free Subspaces
  1. 1.
    The first criterion is the general Lidar-Chuang-Whaley theorem [32]. For a DFS $\mathcal{H}_{\text{DFS}}=\text{span}\left[\{\lvert\psi_{k}\rangle:\ 1\leq k\leq m\}\right]$, all basis states must be degenerate eigenstates of all the jump operators $\hat{L}_{i}$,
@@
    For a schematic of this we show in Fig. 1 that the jump operator $\hat{L}_{i}$ does not couple to states outside of the DFS. At the same time it can couple a generic state in the Hilbert space to a state inside of the DFS. Note this is a feature which only non-Hermitian operators can have.
  2. 2.
    The DFS $\mathcal{H}_{\text{DFS}}$ must also be invariant under the effective Hamiltonian
@@ III Model
Elimination of the excited states
We work in the highly-detuned, weakly-driven limit where the detuning $\Delta_{e}=(\omega_{1}+\omega_{2})/2-\omega_{e}$ is much larger than the coupling rates, $\lvert\Delta_{e}\rvert\gg g,\lvert\Omega_{1}\rvert,\lvert\Omega_{2}\rvert$. This allows us to eliminate the excited states $\lvert i_{e}\rangle$, $i\in\{-1,0,1\}$. We also assume that the cavity coupling is much larger than the drive, $g\gg\lvert\Omega_{1}\rvert,\lvert\Omega_{2}\rvert$, in order to eliminate Raman coupling between |-$\lvert\text{-}1_{g}\rangle$ and $\lvert 1_{g}\rangle$.
Elimination of the cavity
In addition, we assume that the lifetime of a cavity photon is much shorter than the typical timescale of dynamics within the atomic ground states. This is based on $\kappa\gg|\sqrt{N}g\Omega_{1}/\Delta_{e}|,\ |\sqrt{N}g\Omega_{2}/\Delta_{e}|$, where the $\sqrt{N}$ scaling comes from the collective enhancement of the atom-cavity coupling. To eliminate the cavity degrees of freedom we also need them to follow the dynamically modulated drives $\kappa\gg\partial_{t}\eta/\eta,\partial_{t}\Omega_{1}/\Omega_{1},\partial_{t}\Omega_{2}/\Omega_{2}$ with $\partial_{t}f=\partial f/\partial t$.
@@
where we introduced the dressed cavity detuning $\Delta_{c}^{\prime}=(\omega_{1}+\omega_{2})/2-\omega_{c}-Ng^{2}/\Delta_{e}$.
The jump operator are superpositions of the collective lowering and raising operators,
$\hat{J}_{-}=\sum_{j=1}^{N}\lvert-1_{g}\rangle_{j}\langle 0_{g}\rvert_{j}+\lvert 0_{g}\rangle_{j}\langle 1_{g}\rvert_{j}=\hat{J}_{+}^{\dagger}$ (9)
which is generated by cavity decay inducing collective emission which takes an atom from $\lvert 1_{g}\rangle\rightarrow\lvert 0_{g}\rangle$ or $\lvert 0_{g}\rangle\rightarrow\lvert\text{-}1_{g}\rangle$. While this process is in general collectively enhanced, we have introduced the single-particle decay rate
@@ IV Engineering a Decoherence-Free Subspace
Therefore, a DFS is given by the span of all states which meet this criterion, $\mathcal{H}_{\text{DFS}}=\mathrm{span}\left[\{\lvert\psi\rangle:\,\hat{L}\lvert\psi\rangle=0\right]$, for given values of $\mu(t)$ and $\chi(t)$. We now follow the methodology of Ref. [27] to find analytical expressions for the DFS states.
@@
We can represent both the states and operators using Schwinger bosons [59, 57]. We define the annihilation (creation) operator $\hat{b}_{i}$ ($\hat{b}_{i}^{\dagger}$) for $i\in\{-1,0,1\}$ which annihilates (creates) an atom in the state $\lvert i_{g}\rangle$, which satisfy the canonical commutation relations $\left[\hat{b}_{i},\hat{b}_{j}^{\dagger}\right]=\delta_{i\,j}$. This allows us to write
@@ IV.1 Jump Operator Eigenstates
where $\lvert 0\rangle$ is the bosonic vacuum state. Here, $\mathcal{N}_{\vec{k}}$ is a normalization factor which is more complicated than the usual product of factorials and is explicitly calculated in Appendix B. The vector $\vec{k}=(k_{1},k_{2},k_{3})$ defines the state where $0\leq k_{i}\leq N$ for $i\in\{1,2,3\}$ such that $k_{1}+k_{2}+k_{3}=N$. As operators $\hat{c}_{i}^{\dagger}$ are not true bosonic creation operators, states $\lvert\psi_{\vec{k}}\rangle$ $\lvert\psi_{\vec{k}^{\prime}}\rangle$ with $\vec{k}^{\prime}\neq\vec{k}$ are in general not orthogonal. Still, the set $\{\lvert\psi_{\vec{k}}\rangle\}$ for all valid $\vec{k}$ is linearly independent and thus forms a basis for the entire Hilbert space with dimension matching Eq. (14).
@@
Before we describe the DFS in greater detail we will briefly discuss the special case $\mu=0$. This is a special case where $\mathbf{L}$ is not diagonalizable, so the above process cannot be applied. Even so, for $\mu=0$ $\mathbf{L}$ generally has several dark states. The most relevant in this work is the previously mentioned initial state
@@
IV.2 DFS Structure
In the previous section we established that the DFS can be classified by an integer number $C=\chi/\left(\sqrt{2}\mu\right)$. Therefore, the DFS defined by a specific $C$ will be spanned by all the eigenstates $\lvert\psi_{\vec{k}}\rangle$ which satisfy $k_{3}-k_{1}=C$. Hence, we may formally define the DFS for a specific choice of $C$ and $\mu$ to be
@@ IV.2 DFS Structure
Let us now examine the structure of these subspaces in more depth.
In Fig. 3(a), we show the structure of each possible DFS for $N=5$ atoms. Although $\vec{k}$ is a three-dimensional object, the normalization condition $k_{1}+k_{2}+k_{3}=N$ means there are only two independent degrees of freedom, so we may choose these to be $k_{1},k_{3}$. Lines of constant $k_{3}-k_{1}$ are shown connecting degenerate states together. Consider, for example, the orange line going through $(k_{1},k_{3})=(0,3)$ and $(1,4)$. This represents $\mathcal{H}_{\text{DFS}}^{(3)}(\mu)=\mathrm{span}\left[\left\{\lvert\psi_{0,2,3}\rangle,\lvert\psi_{1,0,4}\rangle\right\}\right]$ and corresponds to $C=3$ and the drive $\chi=3\sqrt{2}\mu$. There are only two eigenstates in this subspace because there are no other $\vec{k}$ which satisfy $k_{3}-k_{1}=3$ and $k_{1}+k_{2}+k_{3}=5$. In general, we find that the dark states $\lvert\psi_{\vec{k}}\rangle$ for a specific $C$ satisfy
@@ IV.3 Complementary Subspace
When working with non-orthogonal basis states, it is challenging to find the orthogonal complement of a subspace. As we have discussed in Sec. II, this is particularly relevant for calculating the orthogonal complement of the DFS $\mathcal{H}_{\text{DFS}}^{(C)}$ given in Eq. (27). Fortunately, by defining the operators $\hat{c}_{i}$ and $\hat{d}_{i}$ we have very natural access to the complementary Hilbert space which is based on the relation
@@
Using this definition and property we can write down the orthogonal complement of $\mathcal{H}_{\text{DFS}}^{(C)}$ as
@@ V Time Evolution
We now consider dynamical control of $\mu$ and $\chi$ which is realized by changing the drive amplitude $\eta$ and the ratio of $\Omega_{2}/\Omega_{1}$. Consequently, the DFS $\mathcal{H}_{\text{DFS}}^{(C)}(\mu(t))$ will change while we fix the integer $C=\chi/\left(\sqrt{2}\mu\right)$. By keeping $\Omega_{1}$ constant, the cavity-induced spontaneous emission rate $\Gamma_{c}$ remains constant. We initialize the system at $t=0$ in the ground state $\lvert\psi(0)\rangle$ [(see Eq. (26)] with $\mu=0=\chi$. We then vary $\Omega_{2}(t)$ and $\eta(t)$ until at final time $t_{f}$ we reach $\Omega_{2}(t_{f})=\Omega_{1}$, meaning that $\mu(t_{f})=1$ and $\chi(t_{f})=\sqrt{2}C$. We target $\mu(t_{f})=1$ so that $\hat{L}$ is Hermitian, whereupon we find its eigenstates are highly entangled (see Appendix C). However, we note that one may choose any $\mu\geq 0$ as the target DFS and our formalism will still hold.
@@ V.2 Adiabatic Ramp
At $t_{f}=1/\beta$, we reach $\mu(t_{f})=\beta t_{f}=1$, so the Hamiltonian and jump operator of the system have reached their final values and the system should be ideally in some pure state within $\mathcal{H}_{\text{DFS}}^{(C)}(1)$. This ideal result will never hold perfectly because the dynamics are not perfectly adiabatic unless the evolution takes infinite time, meaning that as $\mu$ changes, the changing Hamiltonian and jump operator will cause some component of the state $\hat{\rho}_{\mathrm{at}}(t)$ to enter the orthogonal complement $\mathcal{H}_{\text{CS}}$. To achieve a final state which meets the threshold for purity and target overlap as given in Eq. (35), we numerically calculate the largest $\beta$ where the threshold is satisfied, as this gives the smallest still sufficient ramping time $t_{f}=1/\beta$.
@@ V.3 Adiabatic Shortcut
The first question that arises is to which states one can shortcut with such a Hamiltonian. It can be proven that we can only shortcut to states that do not possess interparticle entanglement. This can be seen by matching the possible terms in Eq. (37) to the terms which appear in $\langle\psi_{\vec{n}}^{\perp}\rvert\partial_{t}\psi_{\vec{k}}\rangle$ [see Eq. (91) in the Appendix], but the proof does not appear in this work. The intuitive explanation for this finding is that we have initialized the state with an unentangled product state $\lvert\psi(0)\rangle$ [see Eq. (26)] and the shortcut Hamiltonian cannot create entanglement because it is made of single-particle terms. This means we can shortcut exactly the states of the form $\mathcal{N}(\hat{c}_{i}^{\dagger})^{N}\lvert 0\rangle$ for $1\leq i\leq 3$ which describe unentangled states.
To be specific in the following, we distinguish shortcut Hamiltonians $\hat{H}_{s}$ that are used for the preparation of $\lvert\psi_{N,0,0}\rangle$ and $\lvert\psi_{0,0,N}\rangle$ which are both in a one-dimensional DFS from a Hamiltonian that shortcuts to $\lvert\psi_{0,N,0}\rangle$ which resides in a DFS of maximum dimension given in Eq. (29) for $C=0$.

### v2

# Engineering tunable decoherence-free subspaces with collective atom-cavity interactions

@@ I Introduction
A decoherence-free subspace (DFS) allows for exactly this, by taking advantage of the structure of quantum decoherence in order to remove it completely in a quantum system [24, 25, 26, 27, 28]. In a DFS, one engineers the dissipation to stabilize the desired quantum states against the dominant external noise sources. Contrastingly, DFSs arise in quantum resetting schemes [29] where this stabilization prohibits the system from returning to its original state. A promising approach in engineering dissipation is the use of dissipative cavity modes, where one controls the coupling of an atomic ensemble to the cavity mode via a dynamical driving protocol [30, 27, 31, 32]. These protocols have the advantage of removing the need for full control over all constituents of the quantum system, while the global coupling to the cavity mode creates a collective DFS. Such implementations have been discussed in the context of quantum computation and quantum sensing [33, 34, 35, 36, 37] where they allow stable coherence over longer timescales. However, the reduction of the required control in such implementations comes at the cost of their universality because, in general, they only protect against a subset of all possible errors. As a result, useful platforms and protocols which formulate a highly controllable DFS are extremely valuable.
@@
In particular, states can be prepared using time dependent dynamics where the state remains in an instantaneous DFS at every point in time. This allows states to be prepared using adiabatic shortcuts [46, 24] employing experimentally relevant driving schemes. Within this system, we show that a transitionless quantum driving method is possible. These methods allow one to prepare a decoherence-free state without any loss of purity and in timescales not limited by any error rate. We contrast this with typical methods of preparing a DFS which require dissipative state preparation [37] or adiabatic change to the system [47, 48, 49], and take orders of magnitude longer than our adiabatic shortcuts. These shorter timescales mean there is far less opportunity to have errors against which the DFS does not protect.
We also provide a set of operators which drive between degenerate states in a decoherence-free qubit without causing any decoherence. This provides a step towards realizing a controllable DFS for quantum simulation [30, 50] and sensing [51, 52, 53, 54, 55] in atom-cavity setups.
This article is organized as follows. Sec. II is an outline of the basic requirements for a subspace to be a stable DFS. In Sec. III we introduce the collective atom-cavity model which will be used to create DFSs. From this, in Sec. IV, we calculate the eigenstates of the decoherence operator to determine the dimension and structure of every DFS in the system. Next, Sec. V is a discussion of various time evolution methods used to prepare a DFS. Sec. VI defines and examines effective Pauli operators that drive within a two-dimensional DFS. We conclude in Sec. VII with a discussion of future work.
@@ II Stable Decoherence-Free Subspaces
(1) The first criterion is the general Lidar-Chuang-Whaley theorem [33]. For a DFS $\mathcal{H}_{\text{DFS}}=\text{span}\left[\{\lvert\psi_{k}\rangle:\ 1\leq k\leq m\}\right]$, all basis states must be degenerate eigenstates of all the jump operators $\hat{L}_{i}$,
@@
For a schematic of this we show in Fig. 1 that the jump operator $\hat{L}_{i}$ does not couple to states outside of the DFS. At the same time it can couple a generic state in the Hilbert space to a state inside of the DFS. Note that this is a feature unique to non-Hermitian operators, as a Hermitian $\hat{L}_{i}$ only causes dephasing which cannot change the probabilities of measuring each eigenstate of $\hat{L}_{i}$.
(2) The DFS $\mathcal{H}_{\text{DFS}}$ must also be invariant under the effective Hamiltonian
@@ III Model
(a) Elimination of the excited states:We work in the highly-detuned, weakly-driven limit where the detuning $\Delta_{e}=(\omega_{1}+\omega_{2})/2-\omega_{e}$ is much larger than the coupling rates, $\lvert\Delta_{e}\rvert\gg g,\lvert\Omega_{1}\rvert,\lvert\Omega_{2}\rvert$. This allows us to eliminate the excited states $\lvert i_{e}\rangle$, $i\in\{-1,0,1\}$, so we are left with only the three ground states and the cavity. We also assume that the cavity coupling is much larger than the drive, $g\gg\lvert\Omega_{1}\rvert,\lvert\Omega_{2}\rvert$, in order to eliminate Raman coupling between |-$\lvert\text{-}1_{g}\rangle$ and $\lvert 1_{g}\rangle$.
(b)Elimination of the cavity: In addition, we assume that the lifetime of a cavity photon is much shorter than the typical timescale of dynamics within the atomic ground states. This is based on $\kappa\gg|\sqrt{N}g\Omega_{1}/\Delta_{e}|,\ |\sqrt{N}g\Omega_{2}/\Delta_{e}|$, where the $\sqrt{N}$ scaling comes from the collective enhancement of the atom-cavity coupling. To eliminate the cavity degrees of freedom we also need them to follow the dynamically modulated drives $\kappa\gg\partial_{t}\eta/\eta,\partial_{t}\Omega_{1}/\Omega_{1},\partial_{t}\Omega_{2}/\Omega_{2}$ with $\partial_{t}f=\partial f/\partial t$.
@@
where we introduced the dressed cavity detuning $\Delta_{c}^{\prime}=(\omega_{1}+\omega_{2})/2-\omega_{c}-Ng^{2}/\Delta_{e}$. We note that this detuning may be set to zero without fundamentally changing the system dynamics.
The jump operators are a superposition of the collective lowering and raising operators,
$\hat{J}_{-}=\sum_{j=1}^{N}\lvert-1_{g}\rangle_{j}\langle 0_{g}\rvert_{j}+\lvert 0_{g}\rangle_{j}\langle 1_{g}\rvert_{j}=\hat{J}_{+}^{\dagger}.$ (9)
These are generated by cavity decay inducing collective emission which takes an atom from $\lvert 1_{g}\rangle$ to $\lvert 0_{g}\rangle$ or $\lvert 0_{g}\rangle$ to |-$\lvert\text{-}1_{g}\rangle$. While this process is in general collectively enhanced, we have introduced the single-particle decay rate
@@ IV Engineering a Decoherence-Free Subspace
Therefore, a DFS is given by the span of all states which meet this criterion, $\mathcal{H}_{\text{DFS}}=\mathrm{span}\left[\{\lvert\psi\rangle:\,\hat{L}\lvert\psi\rangle=0\}\right]$, for given values of $\mu(t)$ and $\chi(t)$. We now follow the methodology of Ref. [27] to find analytical expressions for the DFS states.
@@
We can represent both the states and operators using Schwinger bosons [60, 58]. We define the annihilation (creation) operator $\hat{b}_{i}$ ($\hat{b}_{i}^{\dagger}$) for $i\in\{-1,0,1\}$ which annihilates (creates) an atom in the state $\lvert i_{g}\rangle$, and satisfy the canonical commutation relations $\left[\hat{b}_{i},\hat{b}_{j}^{\dagger}\right]=\delta_{i\,j}$. This allows us to write
@@ IV.1 Jump operator eigenstates
where $\lvert 0\rangle$ is the bosonic vacuum state. Here, $\mathcal{N}_{\vec{k}}$ is a normalization factor which is more complicated than the usual product of factorials and is explicitly calculated in Appendix B. The vector $\vec{k}=(k_{1},k_{2},k_{3})$ defines the state where $0\leq k_{i}\leq N$ for $i\in\{1,2,3\}$ such that $k_{1}+k_{2}+k_{3}=N$. As operators $\hat{c}_{i}^{\dagger}$ are not true bosonic creation operators, states $\lvert\psi_{\vec{k}}\rangle$, $\lvert\psi_{\vec{k}^{\prime}}\rangle$ with $\vec{k}^{\prime}\neq\vec{k}$ are in general not orthogonal. Still, the set $\{\lvert\psi_{\vec{k}}\rangle\}$ for all valid $\vec{k}$ is linearly independent and thus forms a basis for the entire Hilbert space with dimension matching Eq. (14).
@@
There are a few considerations to note.
(a) Non-Integer $C$: In an experiment, $C$ may not be an exact integer due to fluctuations in $\chi$ and $\mu$. In this case the system will not exactly reach a dark state but will approach a mixed state with a large overlap with the DFS and a small overlap with the orthogonal subspace of the DFS. The small overlap will be determined by the magnitude of fluctuations and therefore we can assume it to be negligible provided that the fluctuations are perturbative. Note that for the special case of the cavity-drive $\eta$ being turned off, then $\chi=0$ and $C=0$ is exactly an integer. In this work, we will assume that $C$ is an integer for simplicity.
(b) Zero Collective Pump Ratio: We will now briefly discuss the case of $\mu=0$. This is a special case where $\mathbf{L}$ is not diagonalizable, so the above process cannot be applied. Even so, for $\mu=0$ $\mathbf{L}$ generally has several dark states. The most relevant in this work is the previously mentioned initial state
@@
IV.2 Decoherence-free subspace structure
In the previous section, we established that the DFS can be classified by an integer number $C=\chi/\left(\sqrt{2}\mu\right)$. Therefore, the DFS defined by a specific $C$ will be spanned by all the eigenstates $\lvert\psi_{\vec{k}}\rangle$ which satisfy $k_{3}-k_{1}=C$. Hence, we may formally define the DFS for a specific choice of $C$ and $\mu$ to be
@@ IV.2 Decoherence-free subspace structure
Let us now examine the structure of these subspaces in more depth. In Fig. 3(a), we show the structure of each possible DFS for $N=5$ atoms. Although $\vec{k}$ is a three-dimensional object, the condition $k_{1}+k_{2}+k_{3}=N$ means there are only two independent degrees of freedom, so we may choose these to be $k_{1},k_{3}$. Lines of constant $k_{3}-k_{1}$ are shown connecting degenerate states together. Consider, for example, the orange line going through $(k_{1},k_{3})=(0,3)$ and $(1,4)$. This represents $\mathcal{H}_{\text{DFS}}^{(3)}(\mu)=\mathrm{span}\left[\left\{\lvert\psi_{0,2,3}\rangle,\lvert\psi_{1,0,4}\rangle\right\}\right]$ and corresponds to $C=3$ and the drive $\chi=3\sqrt{2}\mu$. There are only two eigenstates in this subspace because there are no other $\vec{k}$ which satisfy $k_{3}-k_{1}=3$ and $k_{1}+k_{2}+k_{3}=5$. In general, we find that the dark states $\lvert\psi_{\vec{k}}\rangle$ for a specific $C$ satisfy
@@ IV.3 Complementary subspace
When working with non-orthogonal basis states, it is challenging to find the orthogonal complement of a subspace. As we have discussed in Sec. II, this is particularly relevant for calculating the orthogonal complement of the DFS $\mathcal{H}_{\text{DFS}}^{(C)}(\mu)$ given in Eq. (27). Fortunately, by defining the operators $\hat{c}_{i}$ and $\hat{d}_{i}$ we have very natural access to the complementary Hilbert space which is based on the relation
@@
Using this definition and property we can write down the orthogonal complement of $\mathcal{H}_{\text{DFS}}^{(C)}(\mu)$ as
@@ V Time Evolution
We now consider dynamical control of $\mu$ and $\chi$ which is realized by changing the drive amplitude $\eta$ and the ratio of $\Omega_{2}/\Omega_{1}$. Consequently, the DFS $\mathcal{H}_{\text{DFS}}^{(C)}(\mu(t))$ will change while we fix the integer $C=\chi/\left(\sqrt{2}\mu\right)$. By keeping $\Omega_{1}$ constant, the cavity-induced spontaneous emission rate $\Gamma_{c}$ remains constant. We initialize the system at $t=0$ in the ground state $\lvert\psi(0)\rangle$ [see Eq. (26)] with $\mu=0=\chi$. We then vary $\Omega_{2}(t)$ and $\eta(t)$ until at final time $t_{f}$ we reach $\Omega_{2}(t_{f})=\Omega_{1}$, meaning that $\mu(t_{f})=1$ and $\chi(t_{f})=\sqrt{2}C$. We target $\mu(t_{f})=1$ so that $\hat{L}$ is Hermitian, whereupon we find its eigenstates are highly entangled (see Appendix C). However, we note that one may choose any $\mu\geq 0$ as the target DFS and our formalism will still hold.
@@ V.2 Adiabatic ramp
At $t_{f}=1/\beta$, we reach $\mu(t_{f})=\beta t_{f}=1$, so the Hamiltonian and jump operator of the system have reached their final values and the system should be ideally in some pure state within $\mathcal{H}_{\text{DFS}}^{(C)}(1)$. This ideal result will never hold perfectly because the dynamics are not perfectly adiabatic unless the evolution takes infinite time, meaning that as $\mu$ changes, the changing Hamiltonian and jump operator will cause some component of the state $\hat{\rho}_{\mathrm{at}}(t)$ to enter the orthogonal complement $\mathcal{H}_{\text{CS}}(\mu)$. To achieve a final state which meets the threshold for purity and target overlap as given in Eq. (35), we numerically calculate the largest $\beta$ where the threshold is satisfied, as this gives the smallest still sufficient ramping time $t_{f}=1/\beta$.
@@ V.3 Adiabatic shortcut
The first question that arises is to which states one can shortcut with such a Hamiltonian. It can be proven that we can only shortcut to states that do not possess interparticle entanglement. This can be seen by matching the possible terms in Eq. (37) to the terms which appear in $\langle\psi_{\vec{n}}^{\perp}\rvert\partial_{t}\psi_{\vec{k}}\rangle$ [see Eq. (97) in the Appendix], but the proof does not appear in this work. The intuitive explanation for this finding is that we have initialized the state with an unentangled product state $\lvert\psi(0)\rangle$ [see Eq. (26)] and the shortcut Hamiltonian cannot create entanglement because it is made of single-particle terms. This means we can shortcut exactly the states of the form $\mathcal{N}(\hat{c}_{i}^{\dagger})^{N}\lvert 0\rangle$ for $i\in\{1,2,3\}$ which describe unentangled states.
To be specific in the following, we distinguish shortcut Hamiltonians $\hat{H}_{s}$ that are used for the preparation of $\lvert\psi_{N,0,0}\rangle$ and $\lvert\psi_{0,0,N}\rangle$ which are each in a one-dimensional DFS from a Hamiltonian that shortcuts to $\lvert\psi_{0,N,0}\rangle$ which resides in a DFS of maximum dimension given in Eq. (29) for $C=0$.
We note that despite only being able to shortcut directly to unentangled states, the multi-dimensional $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$ contains highly entangled states which are degenerate with $\lvert\psi_{0,N,0}\rangle$. In Sec. VI we derive a driving scheme between states in a two-dimensional DFS which avoids decoherence. For the cases of $N\in\{2,3\}$, $\mathcal{H}_{\text{DFS}}^{(0)}(\mu)$ is two-dimensional, meaning that we can drive between $\lvert\psi_{0,N,0}\rangle$ and its degenerate states without introducing any decoherence.